QUESTION 23

  1. Two alternatives have the following cash flows:

    Alternative

    Year A B
    0 -$10000 -$15000
    1 +$3000 +$4400
    2 +$3000 +$4400
    3 +$3000 +$4400
    4 +$3000 +$4400

    Assuming a 5% MARR, use incremental analysis to determine which alternative should be selected?
Answer #1

incremental initial cost (B-A) = 15000 - 10000 = 5000

Incremental annual benefit = 4400 - 3000 = 1400

Let incremental IRR be I%, then

-5000 + 1400 * (P/A, I%,4) = 0

1400 * (P/A, I%,4) = 5000

Dividing by 200

7 * (P/A, I%,4) = 25

When I = 5%, value of 7 * (P/A, I%,4) = 24.821654

When I = 4%, value of 7 * (P/A, I%,4) = 25.409267

using interpolation

I = 4% + (25.409267 - 25) / (25.409267 - 24.821654) *(5% -4%)

I = 4% + 0.6964%

I = 4.70%

As incremental IRR < MARR, so option A should be selected